A simple circuit is constructed by connecting one voltage source E Volt with one resistor resistor R Ohm and current is measured 1.48 Amps. If the voltage of the source is TWICE and resitor is ONE-THIRD of their previous values, what is the new current?
Answer #1

current i= 1.48 A

using ohms law

E = i R

i= E/R

so 1.48= E/R -------------------------(1)

new voltage = 2 E and new resistance = R / 3

new current = (2 E)/ (R / 3)

new current = 6 (E / R)

we plugin the value of E/R = 1.48 from equation 1

new current= 6 * I = 6 * 1.48

new currenr= 8.88A